OpenStudy (anonymous):

Solve for x. 2(x – 4) + 3 = 2x – 5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Step 1: Simplify both sides of the equation

OpenStudy (anonymous):

A. no solution B. –1 C. 1 D. identity

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2%281+%E2%80%93+4%29+%2B+3+%3D+2x+%E2%80%93+5+find+x its 1 but there is the proof

OpenStudy (anonymous):

2x−5=2x−5

OpenStudy (anonymous):

Step 2: Subtract 2x from both sides. 2x−5−2x=2x−5−2x −5=−5

OpenStudy (anonymous):

I thought so, i was confused. Thanks to both of yall.

OpenStudy (anonymous):

Step 3: Add 5 to both sides. −5+5=−5+5 0=0 Answer: All real numbers are solutions.

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

you welcome @JCGarcia550

OpenStudy (anonymous):

how did they guy asking the question get the medal?

OpenStudy (anonymous):

how bout you give me one ill give you one?

OpenStudy (anonymous):

@mariah_blahhh